OpenStudy (anonymous):

solve fo x if cos(x-30)=sin3x

OpenStudy (anonymous):

you know that sin(90-x)=cos x now solve it!!

OpenStudy (anonymous):

explanation???? cant reall focus at the moment paper 3 test tmrw

OpenStudy (anonymous):

As said before sin(90-x)=cos(x) so here sin(90-3x)=cos(x-30) >>>sin(3x)=sin(90-(x-30)) >>>>sin 3x=sin(120-x) >>>>>3x=120-x so x=30 got it?

OpenStudy (anonymous):

got that awer need general solution

OpenStudy (anonymous):

thanxfor that though